# Introducing sgboost: A Practical Guide and Implementation of sparse-group boosting in R

### v1

@@ 2.3.2 Model coefficients
We see that the effect sizes differ between the two perspectives. The variable X15 for example has a more extreme model coefficient of -0.272 in the aggregated case compared to the coefficient of -0.249 derived only from the group base-learner. Consequently, also the ordering differs. X11 has a greater absolute model coefficient from the group than X13, but in the aggregated version the absolute model coefficient of X13 exceeds the one of X11.
@@ 2.3.3 Plotting model coefficients and importance
With plot_effects() we can plot the effect sizes of the sparse-group boosting model in relation to the relative importance to get an overall picture of the model. Through the parameter ’plot_type’ one can choose the type of visualization. ’radar’ refers to a radar plot using polar coordinates. Here the angle is relative to the cumulative relative importance of predictors and the radius is proportional to the effect size. ’clock’ does the same as ’radar’ but uses clock coordinates instead of polar coordinates. ’scatter’ uses the effect size as the y-coordinate and the cumulative relative importance as the x-axis in a classical Scatter plot.
@@
Figure 3: Effect size and
@@ 3 Real data
In figure 6, we look at the variable importance with the default values, plotting all 27 selected predictors of which 8 are groups, the latter having a relative variable importance of 22 percent. The most important base-learner is the individual variable ’S1.8b’ indicating whether farming journals are being used and the most important group is the social asset group followed by the group consisting of the four considered regions.
@@
4 The sparse-group boosting algorithm
General functional gradient descent Algorithm ([3])
  1. 1.
    Define base-learners of the structure $h:\mathbb{R}^{n\times p}\to\mathbb{R}$
  2. 2.
    Initialize $m=0$ and $\widehat{f}^{[0]}\equiv 0$ or $\widehat{f}^{[0]}\equiv\overline{y}$
  3. 3.
    Set $m=m+1$ and compute the negative gradient $\frac{\partial}{\partial f}l(y,f)$ and evaluate it at $\widehat{f}^{[m-1]}$. Doing this yields the pseudo-residuals $u_{1},...,u_{n}$ with
    $u_{i}^{[m]}=\frac{\partial}{\partial f}l(y_{i},f)|_{f=\widehat{f}^{[m-1]}},$
    for all $i=1,...,n$
  4. 4.
    Fit the base-learner $h$ with the response $(u_{1}^{[m]},...,u_{n}^{[m]})$ to the data. This yields $\widehat{h}^{[m]}$, which is an approximation of the negative gradient
  5. 5.
    Update
    $\widehat{f}^{[m]}=\widehat{f}^{[m-1]}+\eta\cdot\widehat{h}^{[m]}$
    here $\eta$ can be seen as learning rate with $\eta\in]0,1[$
  6. 6.
    Repeat steps 3,4 and 5 until $m=M$
‘sgboost‘ uses the following adjustment to the general functional gradient descent algorithm. The sparse-group-boosting optimization problem is formulated as the sum of mean squared error (linear regression) or negative log-likelihood (generalized linear models e.g. logistic regression) within each boosting iteration in steps 3 and 4. The base-learner with the smallest evaluated loss function will be selected.
Linear regression:
  • Group base-learner:
    $\operatorname*{arg\,min}_{g\leq G}\min_{\beta^{(g)}\in\mathbb{R}^{p}}\left(\frac{1}{2n}\rVert u^{[m]}-X^{(g)}\beta^{(g)}\rVert_{2}^{2}+\lambda^{(g)}\rVert\beta^{(g)}\rVert_{2}^{2}\right).$
  • Individual base-learner:
    $\operatorname*{arg\,min}_{g\leq G,j\leq p_{g}}\min_{\beta^{(g)}_{j}\in\mathbb{R}}\left(\frac{1}{2n}\rVert u^{[m]}-X^{(g)}_{j}\beta^{(g)}_{j}\rVert_{2}^{2}+\lambda^{(g)}_{j}\rVert\beta^{(g)}_{j}\rVert_{2}^{2}\right).$
If the group base-learner with the smallest loss as defined above has a smaller loss than the individual base-learner with the smallest loss then the group base-learner will be selected in the update and vice versa.
Generalized linear models:
  • Group base-learners:
    $\operatorname*{arg\,min}_{g\in G}\min_{\beta^{(g)}\in\mathbb{R}}\left(-\text{loglik}(u^{[m]},X^{(g)},\beta^{(g)})+\lambda^{(g)}\rVert\beta^{(g)}\rVert_{2}^{2}\right).$
  • Individual base-learners:
    $\operatorname*{arg\,min}_{g\in G,j\in p_{g}}\min_{\beta^{(g)}_{j}\in\mathbb{R}}\left(-\text{loglik}(u^{[m]},X^{(g)}_{j},\beta^{(g)}_{j})+\lambda^{(g)}_{j}\rVert\beta^{(g)}_{j}\rVert_{2}^{2}\right).$
where
  • $df(\lambda^{(g)}_{j})=\alpha$, $df(\lambda^{(g)})=1-\alpha$
  • $df(\lambda)=tr(2H_{\lambda}-H_{\lambda}^{T}H_{\lambda})$ and $H_{\lambda}$ is the Ridge Hat matrix of a base-learner
  • $X^{(g)}$ is the submatrix of $X$ with columns corresponding to group $g$ and $X^{(g)}_{j}$ the $j$-th column of group $g$.
  • $\beta^{(g)}$ and $\beta^{(g)}_{j}$ are the corresponding parameter vectors of the features in group $g$.
  • $\alpha$ adjusts the weight between group and individual base-learners.
  • $\lambda^{(g)},\lambda^{(g)}_{j}$ Ridge penalty parameters.
  • $u^{[m]}$ is the current negative gradient vector from the previous boosting iteration.

### v2

@@ 3.3.2 Model coefficients
We see that the effect sizes differ between the two perspectives. The variable X15, for example, has a more extreme model coefficient of -0.272 in the aggregated case compared to the coefficient of -0.249 derived only from the group base-learner. Consequently, the ordering also differs. X11 has a greater absolute model coefficient from the group than X13, but in the aggregated version, the absolute model coefficient of X13 exceeds the one of X11.
@@ 3.3.3 Plotting model coefficients and importance
With plot_effects() we can plot the effect sizes of the sparse-group boosting model in relation to the relative importance to get an overall picture of the model. Through the parameter ’plot_type’ one can choose the type of visualization. ’radar’ refers to a radar plot using polar coordinates. Here, the angle is relative to the cumulative relative importance of predictors, and the radius is proportional to the effect size. ’clock’ does the same as ’radar’ but uses clock coordinates instead of polar coordinates. ’scatter’ uses the effect size as the y-coordinate and the cumulative relative importance as the x-axis in a classical Scatter plot.
@@
Figure 3: Three visualizations of Effect size vs. relative importance of individual and group base-learners
@@ 3.4 Real data
In figure 6, we look at the variable importance with the default values, plotting all 27 selected predictors of which 8 are groups, the latter having a relative variable importance of 22 percent. The most important base-learner is the individual variable ’S1.8b’, indicating whether farming journals are being used and the most important group is the social asset group, followed by the group consisting of the four considered regions.
@@
3.5 Group selection bias and balance
The function balance() returns the optimal degrees of freedom for each baselearner, such that all groups have equal selection chances if the outcome is not associated with any group. This is especially important in genetic research as group sizes based on genes may vary strongly. Not controlling for group sizes may lead to strong false detection because of a bias towards more complex groups eg. genes. To illustrate this problem, four scenarios are considered. The first three scenarios have three groups, where the first group is a categorical predictor with three categories, the second group is a categorical predictor with two categories, and the third group is a numerical variable simulated with a standard normal distribution. The fourth scenario consists of two groups, the first having 46 members and the second having 4. In the first scenario, the sample size is 50; in the second and third scenarios, the sample size is 500; in the fourth scenario, the sample size is 30, leading to $p>n$. In scenario 3 the outcome variable is i.i.d gamma distributed with shape one and rate 1, Scenarios one, two, and four have standard normal outcomes. Table 1 shows the selection frequencies in these scenarios for each group, for group boosting with three versions of group boosting: one with equal penalties ($\lambda=0.1$), one with equal degrees of freedom df($\lambda$) = 0.5, and one with the degrees of freedom based on the balance() function which implements Algorithm 2. The default settings of 3000 repetitions of i.i.d standard normal outcomes, 20 iterations, a learning rate of 0.5, and a reduction factor of 0.9.
The results suggest, that generally ridge regression with equal penalties leads to the greatest group imbalance, and equal degrees of freedom lead to relatively lower imbalances, especially when two categorical predictors are compared, and the only group adjustment-based group boosting balances the the selection frequencies in all scenarios. In scenario 4, equal lambda only selects the larger group, and in equal degrees of freedom, the chance of the larger group being selected is 2.9 times higher than the smaller one. equal lambda is better at balancing a binary variable with a numerical variable compared to equal df, where equal df is better at balancing group sizes compared to equal lambda. The distribution of the outcome variable seems to not play a great role as the results in scenarios two and three are quite comparable for all three models, which makes the balancing algorithm robust, even if the distribution of the error is not known. Comparing scenario one with two, the balancing algorithm works better with a greater sample size.
Table 1: Selection frequencies in group boosting with degrees of freedom adjustment (group adjustment) compared to ridge regression with equal degrees of freedom (equal df) and equal penalty term (equal lambda). The degrees of freedom used as group adjustment are shown in brackets.
Scenario group equal lambda equal df group adjustment
1 1 0.699 0.453 0.345 (df=0.377)
1 2 0.157 0.364 0.341 (df=0.406)
1 3 0.144 0.183 0.314 (df=0.717)
2 1 0.701 0.407 0.337 (df=0.397)
2 2 0.15 0.419 0.339 (df=0.352)
2 3 0.149 0.174 0.324 (df=0.751)
3 1 0.695 0.417 0.338 (df=0.397)
3 2 0.155 0.408 0.326 (df=0.352)
3 3 0.15 0.175 0.336 (df=0.751)
4 1 1 0.744 0.518 (df=0.394)
4 2 0 0.256 0.482 (df=0.606)
4 Discussion
4.1 Sparse-group boosting
’sgboost’ when applied to high-dimensional grouped data such as ecological data on climate adaptation, reveals meaningful patterns, such as socio-economic and biophysical variables, thereby providing actionable insights for policy and practice. Moreover, the integration of comprehensive visualization tools—such as variable importance plots, coefficient paths, and effect size charts—enhances the interpretability of the models, making it easier for practitioners to understand the contributions of different predictors.
4.2 Group bias
Attempts to reduce selection bias in boosting have been made through equal degrees of freedom using the definition $\text{df}(\lambda)=\text{tr}(2H^{\lambda}-(H^{\lambda})^{2})$ [7] The results of our simulations highlight a fundamental issue in group boosting: a systematic selection bias towards more complex base-learners. This phenomenon follows the principle that "whoever shouts the loudest is rarely right," meaning that larger or more flexible groups are favored in the selection process. The proposed group balancing function balance() in sgboost implementing the group balancing algorithm effectively eliminating this bias by using a simulation-based approach to equalize selection probabilities across different groups.
While this issue is particularly evident in group boosting, it is not limited to this setting. A similar bias can arise in standard boosting when base-learners differ significantly in scale or distribution. This occurs, for example, when comparing binary and numerical variables or in the context of functional regression. The proposed adjustment method provides a systematic way to address these imbalances across various modeling settings and is robust against small sample sizes and varying outcome variable distributions. Even if one assumes a wrong error distribution in the simulation e.g. standard normal, the group adjustment seems to still balance the group selection frequencies if the actual error distribution differs, e.g. gamma errors.
4.3 Limitations of the group balancing algorithm
Despite its benefits, the balancing approach has several challenges: The resampling and iterative adjustments make the method time-intensive. Different combinations of degrees of freedom can yield similar selection frequencies. One way to address this is by fixing the degrees of freedom of a reference base-learner and adjusting only the others. If the learning rate is too high or the number of resamples is too low, the algorithm may fail to converge. To mitigate this, a reduction factor is applied. The method can overshoot adjustments, leading to situations where the ridge regression problem becomes ill-posed (e.g., degrees of freedom approaching zero or leading to negative. This is controlled via predefined bounds (max_df and min_df) which can be used instead of the too-extreme solution.
While the algorithm increases computational cost, it is important to compare this to alternative approaches. Standard group boosting with equal penalties often requires extensive tuning, including thousands of boosting iterations and 25-fold cross-validation to determine optimal stopping. In contrast, the balancing approach achieves the same computational efficiency when considering this tuning overhead. Furthermore, parallelization can significantly reduce runtime. Additionally, equal degrees of freedom approaches also require multiple $\lambda$ values to equalize the selection behavior, making the balancing method computationally more efficient in comparison if the $\lambda$ values are optimized and not the degrees of freedom.